The wave function of a standing wave is defined as y (x,t)=(4.44mm)sin {(0.0340rad/cm)}×sin {(754rad/s)t}

For two travelling waves that make up this standing wave.

Find:

(a) The amplitude?

(b) wavelength?

(c) Frequency?

(d) Wavespeed?

(e) Wave functions?

(f) From the given information, can you determine which harmonic this is? Explain your answer.


1
Expert's answer
2021-04-05T11:07:08-0400




Answers




"\bigstar"

a. amplitude = "\boxed{4.44\over 2}"

= 2.22mm





"\bigstar"

b. "\boxed{ V = f \lambda}"


"\lambda" = "{754\over 0.340 }\times{ 1\over 100}" "\times" "2\pi \over 754"


"\boxed{\lambda}" = "\boxed{2π\over 34}"






"\bigstar"

c. "\boxed{\omega = 2\pi{f}}" = 754

"\implies"

"\boxed f" = "\boxed{754\over 2\pi}"






"\bigstar"

d. "\boxed V" = "\boxed{\omega \over\kappa}"

V = "754\over 0.340" "cm\/s"


"\boxed V" = "\boxed{{754\over 0.340 }\times{ 1\over 100}}" m/s







"\bigstar"

e.

We use "\boxed {2sinAsinB = cos(A-B)-cos(A+B)}"

On comparing we got

"\boxed {2.22[2sin(0.0340x)cos(754t)]}"


On solving we got

First function- "\boxed {2.22cos(0.0340x-754t)}"



second function "\boxed {- [2.22cos(0.0340x+754t)]}"





"\bigstar"

f.

No"\implies" length of the function is not given